# Exploring the statistical properties of the neutron-deficient 109In isotope with the Oslo method

### v1

@@ III.3 Theoretical model calculations
Thus, within the REOM2/RQTBA model, we obtain a qualitative change of the pattern of the low-energy IV $E1$ strength In contrast to neutron-rich nuclei, in the neutron-deficient species, the lowest dipole states are largely driven by proton transitions. Although the neutron peak at the surface, often associated with the neutron skin oscillation, is still notably expressed in the radial behavior of the transition densities, its surface contribution to the strength is compensated by the oscillation in the interior. The 110Sn nucleus thus represents the onset of this qualitative change within the tin isotopic chain.
@@ IV Neutron- and proton-capture cross sections and rates
As was mentioned in the introduction, the NLDs and GSFs extracted with the Oslo method can provide important constraints for calculations of radiative neutron- and proton-capture rates in astrophysical $p$, $s$, $i$, and $r$ processes within the Hauser-Feschbach statistical framework [95]. To address the impact of our new experimental data for 109In, we performed calculations of cross sections and rates of the radiative neutron- and proton-capture reactions, 108In($n,\gamma$)109In and 108Cd($p,\gamma$)109In, respectively, using the talys software. The experimental Oslo method NLD and GSF, combined with the ($\gamma,xn$) data for 115In at energies above the neutron separation energy, were used to produce tabulated NLD and $E1$ strength inputs for talys calculations. In addition to the uncertainties of all normalization parameters, the uncertainty in the $M1$ response and, therefore, the decomposition of the experimental strength into the $E1$ and $M1$ components was taken into account: all $M1$ models available in talys, as well as the $M1$ resonance extracted from the Sn systematics (see Fig. 9), were included in the uncertainty band. Moreover, the optical potentials from both the phenomenological model of Koning and Delaroche [96] and the semi-microscopic Jeukenne-Lejeune-Mahaux model renormalized by the Bruyères-le-Châtel group [97] were included in the experimental uncertainties for the cross sections. The results of the calculations with the experimental Oslo method inputs are shown as shaded yellow bands in Figs. 13 and 14 for neutron and proton capture reactions, respectively. The talys model uncertainty in both cases was also estimated by varying all available NLD, optical model potential, $E1$ and $M1$ strength options. When considering different model combinations, the Skyrme-HFB-QRPA [77] and Gogny-HFB-QRPA [79] $M1$ strengths were employed together with the corresponding $E1$ model calculations. The resulting range of talys model uncertainties is displayed by broad beige bands in Figs. 13 and 14. As default cross sections and rates, also shown in the figures by dashed-dotted lines, the talys software utilizes the CT plus Fermi gas NLD model, the Simplified Modified Lorentzian (SMLO) models for the $E1$ and $M1$ strengths, and the global optical model potential by Koning and Delaroche (see Ref. [68] for more details regarding the used models).
@@
Figure 14: The experimentally constrained cross section (a) and rate (b) of the 108Cd($p,\gamma$)109In reaction, (orange band) shown together with the talys model uncertainty (beige band) and default calculations (dash-dotted line). The experimental data by Skakun et al. [100], Gyürky et al. [101] and Olivas-Gomez et al. [102] are shown in the upper panel. Predictions from the JINA REACLIB [98] and BRUSLIB [99] libraries are displayed in the bottom panel with short-dashed and long-dashed lines, respectively.
@@
The radiative proton capture on stable 108Cd has been previously studied in Refs. [100, 101, 102], and the comparison of the extracted ($p,\gamma$) cross sections with the present results is shown in Fig. 14(a). The work by Olivas-Gomez et al. [102] specifically focuses on constraining the NLD and the GSF models with experimental proton capture cross sections on 108,110Cd, aiming to improve the description of the ($\gamma,n$) and $(\gamma,p)$ rates for the potential $p$-process branching point  [^1] In. Within the uncertainties, our data are fully consistent with the more recent analyses by Gyürky et al. [101] and Olivas-Gomez et al. [102] over a wide range of energies from approximately 3 MeV to 7 MeV and with the older data by Skakun et al. [100] for 7 MeV and higher. This serves as another benchmark for the Oslo method inputs, supporting the assumptions made in the normalization procedure. Similarly to the neutron-capture case, the talys default calculations tend to slightly underestimate the experimental cross sections. The microscopic Skyrme-HF-BCS plus statistical NLD from Ref. [66] and the Skyrme-HFB-QRPA GSF from Ref. [77] were found to be an optimal combination yielding the best agreement with the Oslo data. The latter model, however, underestimates the IVGDR peak quite notably compared to other models available in talys.
@@
In general, our results support the recommendation made in Ref. [102], proposing the Skyrme-HF-BCS NLD by Goriely et al. [66] as a suitable model for cross section calculations in the vicinity of $A\approx 109$ if no experimental data are available. It is important to note, however, that the NLD extracted in this work deviates from this model already at excitation energies of $\approx 6$ MeV. The GSF based on the Skyrme-HF-BCS-QRPA calculations by Goriely and Khan [75], recommended in the same work, does not yield the best overlap with our data when combined with the Skyrme-HF-BCS NLD model. Including an additional $E1$ enhancement at $\approx 7-8$ MeV, with a strength of about $0.5\%$ of the TRK sum rule as extracted in this work, would improve the agreement between the calculated cross section and the experimental data. It is interesting to note that the $M1$ response, despite a great variability in the $M1$ strength distribution between different models [68], is a rather minor contributor to the uncertainty of both the $(n,\gamma)$ and $(p,\gamma)$ cross sections, compared to the $E1$ strength. At proton energies below 3 MeV, the optical model potential dominates the uncertainty band, whereas at higher energies the cross section becomes considerably more sensitive to the choice of the NLD and GSF.
@@
Finally, the ($p,\gamma$) rates were estimated using the Oslo method inputs; the results of the calculations are shown in Fig. 14(b) together with the talys default rate and model uncertainty, as well as the rates from the JINA REACLIB and BRUSLIB libraries. In this case, the JINA REACLIB rate is more consistent with BRUSLIB, but both of them deviate from the experimentally constrained rate at temperaatures 696-9 GK. The $p$-process sensitivity to the 109In($\gamma,p$)108Cd and 109In($\gamma,n)$108In rates, estimated with the detailed balance theorem from the experimentally constrained 108Cd($p,\gamma$)109In and 108In($n,\gamma)$109In rates, appears to be marginal. This is to be expected, considering that production of the 106Cd and 108Cd isotopes is dominated by photodissociation of heavier Cd isotopes. The impact of the new experimental results is primarily indirect: alongside direct ($p,\gamma$) measurements, the newly extracted NLD and GSF offer a new perspective on the performance of theoretical models as well as on the low-lying $E1$ enhancement and its magnitude in neutron-deficient nuclei in the Sn mass region.
@@ V Conclusions and outlook
The cross-sections and rates of the 108Cd($p,\gamma$)109In and 108In($n,\gamma)^{109}$In reactions were estimated within the Hauser-Feshbach framework using the new experimental NLD and GSF as inputs, thus providing important constraints on the presently available model uncertainties. Our results are in excellent agreement with earlier ($p,\gamma$) cross section measurements. The ($n,\gamma$) reaction rate from the JINA REACLIB library, despite its good performance for the neighboring heavier nuclei, deviates significantly from the BRUSLIB predictions and the experimental result. This point could be relevant for future $p$-process calculations utilizing this library.

### v2

@@ III.3 Theoretical model calculations
Thus, within the REOM2/RQTBA model, we obtain a qualitative change of the pattern of the low-energy IV $E1$ strength. In contrast to neutron-rich nuclei, in the neutron-deficient species, the lowest dipole states are largely driven by proton transitions. Although the neutron peak at the surface, often associated with the neutron skin oscillation, is still notably expressed in the radial behavior of the transition densities, its surface contribution to the strength is compensated by the oscillation in the interior. The 110Sn nucleus thus represents the onset of this qualitative change within the tin isotopic chain.
@@ IV Neutron- and proton-capture cross sections and rates
As was mentioned in the introduction, the NLDs and GSFs extracted with the Oslo method can provide important constraints for calculations of radiative neutron- and proton-capture rates in astrophysical $p$, $s$, $i$, and $r$ processes within the Hauser-Feschbach statistical framework [95]. To address the impact of our new experimental data for 109In, we performed calculations of cross sections and rates of the radiative neutron- and proton-capture reactions, 108In($n,\gamma$)109In and 108Cd($p,\gamma$)109In, respectively, using the talys software. The experimental Oslo method NLD and GSF, combined with the ($\gamma,xn$) data for 115In at energies above the neutron separation energy, were used to produce tabulated NLD and $E1$ strength inputs for talys calculations. In addition to the uncertainties of all normalization parameters, the uncertainty in the $M1$ response and, therefore, the decomposition of the experimental strength into the $E1$ and $M1$ components was taken into account: all $M1$ models available in talys, as well as the $M1$ resonance extracted from the Sn systematics (see Fig. 10), were included in the uncertainty band. Moreover, the optical potentials from both the phenomenological model of Koning and Delaroche [96] and the semi-microscopic Jeukenne-Lejeune-Mahaux model renormalized by the Bruyères-le-Châtel group [97] were included in the experimental uncertainties for the cross sections. The results of the calculations with the experimental Oslo method inputs are shown as shaded yellow bands in Figs. 14 and 15 for neutron and proton capture reactions, respectively. The talys model uncertainty in both cases was also estimated by varying all available NLD, optical model potential, $E1$ and $M1$ strength options. When considering different model combinations, the Skyrme-HFB-QRPA [77] and Gogny-HFB-QRPA [79] $M1$ strengths were employed together with the corresponding $E1$ model calculations. The resulting range of talys model uncertainties is displayed by broad beige bands in Figs. 14 and 15. As default cross sections and rates, also shown in the figures by dashed-dotted lines, the talys software utilizes the CT plus Fermi gas NLD model, the Simple Modified Lorentzian (SMLO) models for the $E1$ and $M1$ strengths, and the global optical model potential by Koning and Delaroche (see Ref. [67] for more details regarding the used models).
@@
Figure 15: The experimentally constrained cross section (a) and rate (b) of the 108Cd($p,\gamma$)109In reaction, (orange band) shown together with the talys model uncertainty (beige band) and default calculations (dash-dotted line). The experimental data by Skakun et al. [100], Gyürky et al. [101], Olivas-Gomez et al. [102], and Saha et al. [103] are shown in the upper panel. Predictions from the JINA REACLIB [98] and BRUSLIB [99] libraries are displayed in the bottom panel with short-dashed and long-dashed lines, respectively.
@@
The radiative proton capture on stable 108Cd has been previously studied and presented in Refs. [100, 101, 102, 103], and the comparison of the extracted ($p,\gamma$) cross sections with the present results is shown in Fig. 15(a). The work by Olivas-Gomez et al. [102] specifically focuses on constraining the NLD and the GSF models with experimental proton capture cross sections on 108,110Cd, aiming to improve the description of the ($\gamma,n$) and $(\gamma,p)$ rates for the potential $p$-process branching point  [^1] In. A similar objective to constrain the nuclear input models was also pursued in the most recent study by Saha et al. [103]. Within the uncertainties, our data are fully consistent with the more recent analyses by Gyürky et al. [101], Olivas-Gomez et al. [102], and Saha et al. [103] over a wide range of energies from approximately 3 MeV to 7 MeV and with the older data by Skakun et al. [100] for 7 MeV and higher. This serves as another benchmark for the Oslo method inputs, supporting the assumptions made in the normalization procedure. Similarly to the neutron-capture case, the talys default calculations tend to slightly underestimate the experimental cross sections. The microscopic Skyrme-HF-BCS plus statistical NLD from Ref. [66] and the Skyrme-HFB-QRPA GSF from Ref. [77] were found to be an optimal combination yielding the best agreement with the Oslo data. The latter model, however, underestimates the IVGDR peak quite notably compared to other models available in talys.
@@
In general, our results support the recommendation made in Ref. [102], proposing the Skyrme-HF-BCS NLD by Demetriou et al. [66] as a suitable model for cross section calculations in the vicinity of $A\approx 109$ if no experimental data are available. It is important to note, however, that the NLD extracted in this work deviates from this model already at excitation energies of $\approx 6$ MeV. The GSF based on the Skyrme-HF-BCS-QRPA calculations by Goriely and Khan [75], recommended in the same work, does not yield the best overlap with our data when combined with the Skyrme-HF-BCS NLD model. Including an additional $E1$ enhancement at $\approx 7-8$ MeV, with a strength of about $0.5\%$ of the TRK sum rule as extracted in this work, would improve the agreement between the calculated cross section and the experimental data. The Skyrme-HFB NLD and the simple Lorentzian GSF were recommended in the recent work by Saha et al. in combination with the newly calculated optical potential for p-nuclei (see Ref. [103] for more details) to reproduce the experimental cross sections. In the present work, however, these models are shown to exhibit significant deviations from the Oslo method NLD and GSF. It is interesting to note that the $M1$ response, despite a great variability in the $M1$ strength distribution between different models [67], is a rather minor contributor to the uncertainty of both the $(n,\gamma)$ and $(p,\gamma)$ cross sections, compared to the $E1$ strength. At proton energies below 3 MeV, the optical model potential dominates the uncertainty band, whereas at higher energies the cross section becomes considerably more sensitive to the choice of the NLD and GSF.
@@
Finally, the ($p,\gamma$) rates were estimated using the Oslo method inputs; the results of the calculations are shown in Fig. 15(b) together with the talys default rate and model uncertainty, as well as the rates from the JINA REACLIB and BRUSLIB libraries. In this case, the JINA REACLIB rate is more consistent with BRUSLIB, but both of them deviate from the experimentally constrained rate at temperatures of 696-9 GK. The $p$-process sensitivity to the 109In($\gamma,p$)108Cd and 109In($\gamma,n)$108In rates, estimated with the detailed balance theorem from the experimentally constrained 108Cd($p,\gamma$)109In and 108In($n,\gamma)$109In rates, appears to be marginal. This is to be expected, considering that production of the 106Cd and 108Cd isotopes is dominated by photodissociation of heavier Cd isotopes. The results of the recent study by Saha et al. and the earlier work by Olivas-Gomez et al. highlight a persistent challenge in statistical-model calculations of reaction rates for the p process, namely that direct measurements of ($p,\gamma$) cross sections alone do not fully constrain NLD and GSF models for p-process nuclei. The impact of the new experimental results is therefore primarily indirect: in combination with direct ($p,\gamma$) measurements, the newly extracted NLD and GSF offer a complementary perspective on the performance of theoretical models and on reducing the associated parameter uncertainties, as well as on the low-lying $E1$ enhancement and its magnitude in neutron-deficient nuclei in the Sn mass region.
@@ V Conclusions and outlook
The cross-sections and rates of the 108Cd($p,\gamma$)109In and 108In($n,\gamma)$109In reactions were estimated within the Hauser-Feshbach framework using the new experimental NLD and GSF as inputs, thus providing important constraints on the presently available model uncertainties. Our results are in excellent agreement with earlier ($p,\gamma$) cross section measurements. The ($n,\gamma$) reaction rate from the JINA REACLIB library, despite its good performance for the neighboring heavier nuclei, deviates significantly from the BRUSLIB predictions and the experimental result. This point could be relevant for future $p$-process calculations utilizing this library.